# Preprint Notice

### v3

@@ VII Acknowledgment
Chin Fang wishes to thank his colleague Riccardo Veraldi for valuable assistance with TCP congestion control and kernel optimization in 2024. While Chin was building the data movement appliances subsection III-C, Riccardo’s guidance contributed to the success of this work.
@@
Chin Fang also wishes to acknowledge two Zettar colleagues, Igor Solovyov and Alexander Nazarenko, for their critical engineering contributions to the zx software, which proved instrumental in the work presented herein. Mr. Nazarenko, paired with Chin Fang, earned the Overall Winner title at the SCA19 Data Mover Challenge [40].

### v4

@@ 7. Acknowledgment
Chin Fang wishes to thank his colleague Riccardo Veraldi for valuable assistance with TCP congestion control and kernel optimization in 2024. While Chin was building the data movement appliances (Section s33), Riccardo’s guidance contributed to the success of this work.
@@
Chin Fang also wishes to acknowledge two Zettar colleagues, Igor Solovyov and Alexander Nazarenko, for their critical engineering contributions to the zx software, which proved instrumental in the work presented herein. Mr. Nazarenko, paired with Chin Fang, earned the Overall Winner title at the SCA19 Data Mover Challenge (NSCC, 2019).